# New Construction of Locally Perfect Nonlinear Functions with Application to Sequences Sets with Low Ambiguity Zone

### v2

@@ Abstract
Low Ambiguity Zone (LAZ) sequences play a pivotal role in modern integrated sensing and communication (ISAC) systems. Recently, Wang et al.[1] proposed a definition of locally perfect nonlinear functions (LPNFs) and constructed three classes of both periodic and aperiodic LAZ sequence sets with flexible parameters by applying such functions and interleaving method. Some of these LAZ sequence sets are asymptotically optimal with respect to the Ye-Zhou-Liu-Fan-Lei-Tang bounds under certain conditions. In this paper, we proceed with the constructions of the three new classes of LPNFs with new parameters. By using these LPNFs, we also present a series of LAZ sequence sets with more flexible parameters, addressing the limitations of existing parameter choices. Furthermore, our results show that one of these classes is asymptotically optimal in both the periodic and aperiodic cases, respectively.
@@
Low Ambiguity Zone (LAZ), Locally Perfect Nonlinear Function (LPNF), LAZ Sequence set.
@@ I Introduction
To design sequence sets with various capabilities is one of the significant topics in the research field of communication and radar systems. A critical challenge now lies in the design of sequences that exhibit desirable correlation properties not only in the time domain but also across phase shifts, as effectively captured by the ambiguity function (AF). The AF is a valuable criterion for evaluating the communication and sensing capabilities of sequences[2], sequences exhibiting zero or low ambiguity function sidelobes are essential for mitigating the Doppler effect in advanced Integrated Sensing and Communication (ISAC) systems, enabling accurate signal detection and processing in highly dynamic environments[3]. In 2013, Ding and Feng et al.[4] investigated the theoretical lower bound of the maximum sidelobe of the ambiguity function (AF) based on the Welch bound. Additionally, they presented sequences possessing favorable AF characteristics. Notably, a single sequence among these achieves the desired theoretical lower bound across the entire delay-Doppler (DD) region. Moreover, Wang and Gong et al.[5, 6, 7] proposed several classes of sequences and sequence sets with low AF magnitudes in the entire DD region, leveraging the excellent algebraic properties of additive and multiplicative characters over finite fields. In addition to constructing sequences with low AF sidelobes based on finite field theory, researchers have also attempted to design sequences through optimization algorithms[8, 9, 10, 11]. However, some of these sequences exhibit low AF sidelobes only in specific DD regions.
@@
In fact, for many practically interesting applications in ISAC system, the Doppler frequency range is typically much narrower than the bandwidth of the transmitted signal. Therefore, it’s likely that there’s no need to take into account the entire duration of the signal. In other words, in the study of different applications, it is usually sufficient to focus on the AF sidelobes in the region relevant to specific application requirements. Based on this idea, in the past five years, people have made breakthroughs in the research of AF. In reference [12], Ye et al. first introduced a new definition regarding the low AF sidelobes in a specific region, termed the low/zero ambiguity zone (LAZ/ZAZ). Therefore, they referred to the sequence sets that meet the requirements of the LAZ/ZAZ as LAZ/ZAZ sequence sets, and derived the theoretical lower bounds for the periodic and aperiodic AF sidelobes of unimodular sequence sets in the LAZ, respectively. These bounds are referred to as “Ye-Zhou-Liu-Fan-Lei-Tang bounds” in reference [1], for brevity, we use the name “YZLFLT bounds” in this paper. By applying cubic sequences, certain quadratic phase sequences and cyclic difference set, they also presented four types of the most suitable constructions of sequence family, and these sequences achieve the lower bounds. Cao et al.[13] subsequently identified that a binary sequence set proposed in [14] possesses low ambiguity characteristics over a large area. Note that these sequences are interleaved sequences. Thereafter, Tian et al.[15] defined a set of ZAZ/LAZ sequences based on the perfect nonlinear functions (PNFs) and a new class of mapping functions. Very recently, Wang et al.[1] generalized the concept of PNFs, introducing locally perfect nonlinear functions (LPNFs) and presented a construction of LPNFs. And then, by directly applying these LPNFs and interleaving method, they gave a series of LAZ sequence sets with flexible parameters. Such sequence sets with selecting appropriate parameters are asymptotically optimal with respect to YZLFLT bounds both periodic and aperiodic cases, respectively. Consequently, this drives us to develop sequences with zero or low AF sidelobes characteristics in desired regions. The specific parameters of the above mentioned sequences are listed in Table I. In column 6 of Table I, P.O, P.AO, and AP.AO mean optimal with periodic, asymptotically optimal with periodic, and asymptotically optimal with aperiodic, respectively.
@@
  Ref. Length Set size $Z_{x}$ $Z_{y}$ Optimality   Constraints
  [4] $q-1$ $1$ $q-1$ $q-1$ P.O. $q=p^{l}$, $p$ is a prime.
[12] $p$ $1$ $p$ $p$ P.O. $p$ is an odd prime.
[12] $L$ $1$ $\frac{L}{r}$ $r$ P.O. $\gcd(a,L)=1\,\textrm{if $L$ is odd },$ $r=\gcd(2a,L)>1$.
[12] $L$ $N$ $\frac{\lfloor L/N\rfloor}{r}$ $r$ P.O, if $N|L$. $\gcd(a,L)=1\,\textrm{if L is odd },$ $r=\gcd(2a,L)>1$.
[15] $MN^{2}$ $MN$ $\lfloor\frac{N}{K}\rfloor$ $K$ P.AO. $K<N$, $\gcd(K,N)=1$.
[15] $N(KN+P)$ $N$ $N$ $K$ P.AO. $\gcd(P,NK)=1$.
[15] $p(p-1)$ $p$ $p-1$ $p$ P.AO. $p$ is an odd prime.
[1] $N^{2}$ $N$ $p$ $N$ P.AO, AP.AO. $N$  is an odd number, $p|N(\textrm{the smallest prime})$.
[1] $NK$ $N$ $p$ $K-N+1$ P.AO, AP.AO. $N$  is an odd number, $N<K<2N-1$, $p|N(\textrm{the smallest prime})$.
Theorem5 (i) $NK$ $N$ $p-1$ $K$ P.AO, AP.AO. $K=p^{e},\textrm{$p$\,is an odd prime},$ $N=\varphi(p^{e})$.
Theorem5 (ii) $NK$ $N$ $N$ $p$ P.AO $K=p^{e},\textrm{$p$\,is an odd prime},$ $N=\varphi(p^{e})$.
Theorem6 $NK_{1}$ $N$ $p-1$ $K_{1}-K+2$ P.AO, AP.AO. $K=p^{e},\textrm{$p$\,is an odd prime},$ $K<K_{1}<2K$, $N=\varphi(p^{e})$.
In this paper, we show some generic new constructions on LPNFs in Section 3, and by using these new LPNFs, we present a series of LAZ sequence sets with new parameters. Some sequence sets of the proposed are asymptotically optimal with respect to YZLFLT bounds both periodic and aperiodic cases, respectively. Before doing our main results, in Section 2, we introduce some preliminaries on LAZ sequence sets and LPNFs, including some related results we need in this paper.
@@ II LAZ Sequence Sets and Locally Perfect Nonlinear Functions
In this section we recall some essential concepts concerning LAZ sequence sets, locally perfect nonlinear functions (LPNFs), as well as results that are needed.
@@ II-A LAZ Sequence Sets
In what follows, let $\mathbb{C}$ denote the set of all complex numbers and $\mathcal{S}$ denote the set of sequences of length $D$ which has a size of $M$, all sequences $\mathbf{a}=(a(0),a(1),a(2),\cdots,a(D-1))$ are regarded as unimodular sequences with length $D$. Namely, ${a}(k)\in\mathbb{C}$ and $|{a}(k)|=1$ for all $0\leq k\leq{D-1}$. In particular, $D$ is the period of $\mathbf{a}$ if ${a}(k+D)={a}(k)$ for all $k\geq 0$. $\omega_{N}=e^{\frac{2\pi\sqrt{-1}}{N}}=e^{\frac{2\pi i}{N}}$ is a $N$-th primitive complex root of unity.
Let $\mathbf{a}$ and $\mathbf{b}$ be unimodular sequences with length $D$. Then the aperiodic cross ambiguity function (cross-AF) of $\mathbf{a}$ and $\mathbf{b}$ at time shift $\tau$ and Doppler shift $\nu$ is defined by:
@@
where for $c\in\mathbb{C}$, $\overline{c}$ represents complex conjugation of $c$.
@@
where $t+\tau$ to be taken modulo the period $D$. We denote $\widehat{AF}_{\mathbf{a}}(\tau,\nu)=\widehat{AF}_{\mathbf{a},\mathbf{a}}(\tau,\nu)$ and ${AF}_{\mathbf{a}}(\tau,\nu)={AF}_{\mathbf{a},\mathbf{a}}(\tau,\nu)$, called the auto-AF of $\mathbf{a}$, aperiodic and periodic, respectively.
@@
Theorem 1.
([12], or [1] Lemma 1 and 2) For any unimodular sequence set $\mathcal{S}=(M,D,\Pi,\theta_{\text{max}})$ or $\mathcal{S}=(M,D,\Pi,\hat{\theta}_{\text{max}})$, where $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$. We have
@@
Further assume that $\rho_{\text{LAZ}}=\frac{\theta_{\text{max}}}{\Delta}(\geq 1)$ and $\hat{\rho}_{\text{LAZ}}=\frac{\hat{\theta}_{\text{max}}}{\hat{\Delta}}(\geq 1)$, $\Delta$ and $\hat{\Delta}$ are YZLFLT bounds, $\rho_{\text{LAZ}}$ and $\hat{\rho}_{\text{LAZ}}$ are conventionally referred to as optimality factors. The optimality factors $\rho_{\text{LAZ}}$ and $\hat{\rho}_{\text{LAZ}}$ serve as metrics to evaluate how closely the constructed sequences approach the theoretical lower bounds in both periodic and aperiodic cases, respectively. i.e.
@@ II-B Locally Perfect Nonlinear Functions
Let $K\geq N$ be two positive integers, and $f:$ $\mathbb{Z}_{N}\rightarrow\mathbb{Z}_{K}$ be a function, where $\mathbb{Z}_{N}=\mathbb{Z}/N\mathbb{Z}$. For $0<Z_{x}\leq N$ and $0<Z_{y}\leq K$, The function $f$ is called a locally perfect nonlinear functions and denoted by $(N,K,Z_{x},Z_{y})$-LPNF if for any $-Z_{y}<b<Z_{y}$ and $-Z_{x}<a<Z_{x}$, $a\neq 0$, the equation $f(x+a)-f(x)=b$ has at most one solution $x\in\mathbb{Z}_{N}$. In what follows, to avoid confusion, we denote $(N,K,Z_{x},Z_{y})$-LPNF, as $\langle N,K,Z_{x},Z_{y}\rangle$-LPNF.
@@
If $Z_{y}\geq\lceil\frac{K+1}{2}\rceil$, then $-Z_{y}<b<Z_{y}$ covers any element $b$ in $\mathbb{Z}_{K}$.
The notion of LPNF was introduced in [1] and successfully used to construct LAZ sequence sets by interleaving technique. The result is as follows.
Theorem 2.
([1] Theorem 1) From a $\langle N,K,Z_{x},Z_{y}\rangle$-LPNF we can construct a LAZ sequence set with parameters $(M,D,\Pi,\theta_{\text{max}}\,\,or\,\,\hat{\theta}_{\text{max}})$, where $M=N$, $D=NK$, $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$, $\theta_{\text{max}}=K$, $\hat{\theta}_{\text{max}}=K+Z_{x}-1$.
Wang and Zhou et al.[1] gave a series of LPNFs and constructed a series of LAZ sequence sets with flexible parameters, which are presented in the above theorem. Some of them are asymptotically optimal in both periodic and aperiodic cases. In next section, we will reveal some new constructions of LPNFs, which provides significant improvements over existing frameworks [1]. And then produce a series of LAZ sequence sets with new parameters that are not covered in Theorem 2. Furthermore, a class of the proposed sequence sets demonstrates asymptotically optimal property in both periodic and aperiodic cases.
III New Construction of LPNFs
For $K\geq 3$, let $\mathbb{Z}_{K}^{\ast}$ denote the multiplication group of all invertible elements of the ring $\mathbb{Z}_{K}$. Then $\mathbb{Z}_{K}^{\ast}$ contains $\varphi(K)$ elements, where $\varphi(\cdot)$ is the Euler’s total function. To establish the desired constructions, we need the following lemma.

### v3

# New Constructions of Locally Perfect Nonlinear Functions and Their Application to Sequence Sets With Low Ambiguity Zone

@@ Abstract
Low Ambiguity Zone (LAZ) sequences play a pivotal role in modern integrated sensing and communication (ISAC) systems. Recently, Wang et al. [arXiv:2501.11313] proposed a definition of locally perfect nonlinear functions (LPNFs) and constructed three classes of both periodic and aperiodic LAZ sequence sets with flexible parameters by applying such functions and interleaving techniques. Some of these LAZ sequence sets are asymptotically optimal with respect to the Ye-Zhou-Fan-Liu-Lei-Tang bounds under certain conditions. In this paper, we present constructions of three new classes of LPNFs with new parameters. Based on these LPNFs, we further propose a series of LAZ sequence sets that offer more flexible parameters. Furthermore, our results show that some of these classes are asymptotically optimal in both the periodic and aperiodic cases, respectively.
@@
Ambiguity function (AF), Low Ambiguity Zone (LAZ), Locally Perfect Nonlinear Function (LPNF), Sequence set.
@@ I Introduction
To design sequence sets with various capabilities is one of the significant topics in the research field of communication and radar systems. A critical challenge now lies in the design of sequences that exhibit desirable correlation properties not only in the time domain but also across phase shifts, as effectively captured by the ambiguity function (AF). The AF is a important metric for evaluating the communication and sensing capabilities of sequences[1], sequences exhibiting zero or low AF sidelobes are essential for mitigating the Doppler effect in advanced integrated sensing and communication (ISAC) systems, enabling accurate signal detection and processing in highly dynamic environments[2].
In 2013, Ding and Feng et al.[3] investigated the theoretical lower bound of the maximum sidelobe of the AF based on the Welch bound. Additionally, they presented sequences possessing favorable AF characteristics. Notably, a single sequence among these achieves the desired theoretical lower bound across the entire delay-Doppler (DD) region. Moreover, Wang and Gong et al.[4, 5, 6] proposed several classes of sequences and sequence sets with low AF magnitudes in the entire DD region, leveraging the excellent algebraic properties of additive and multiplicative characters over finite fields. In addition to constructing sequences with low AF sidelobes based on finite field theory, researchers have also attempted to design sequences through optimization algorithms [7, 8, 9, 10]. However, some of these sequences exhibit low AF sidelobes only in specific DD regions.
@@
In fact, for many practically interesting applications in the ISAC system, the Doppler frequency range is typically much narrower than the bandwidth of the transmitted signal [11, 12]. Therefore, it’s likely that there’s no need to take into account the entire duration of the signal. In other words, in the study of different applications, it is usually sufficient to focus on the AF sidelobes in the region relevant to specific application requirements. Based on this idea, in the past five years, people have made breakthroughs in the research of AF. In [13], Ye et al. first introduced a new definition regarding the low AF sidelobes in a specific region, termed the low/zero ambiguity zone (LAZ/ZAZ). Therefore, they referred to the sequence sets that meet the requirements of the LAZ/ZAZ as LAZ/ZAZ sequence sets, and derived the theoretical lower bounds for the periodic and aperiodic AF sidelobes of unimodular sequence sets in the LAZ, respectively. These bounds are collectively known as the Ye–Zhou–Fan–Liu–Lei–Tang bounds. For brevity, we refer to them as the “YZFLLT bounds” throughout this paper. By applying cubic sequences, certain quadratic phase sequences, and cyclic difference sets, they also presented four types of the most suitable constructions of sequence families, and these sequences achieve the lower bounds. Meng et al. [14] recently established a tighter aperiodic bound compared to [28, Theorem 4] under specific LAZ sequence parameters. Subsequently, Cao et al.[15] identified that a binary sequence set proposed in [16] possesses low ambiguity characteristics over a large area. Note that these sequences are interleaved sequences. Thereafter, Tian et al.[17] defined a set of LAZ sequences based on the perfect nonlinear functions (PNFs). Very recently, Wang et al.[18] generalized the concept of PNFs, introducing locally perfect nonlinear functions (LPNFs), and presented three constructions of LPNFs. And then, by directly applying these LPNFs and the interleaving techniques, they gave a series of LAZ sequence sets with flexible parameters. Such sequence sets with selecting appropriate parameters are asymptotically optimal with respect to YZFLLT bounds in both periodic and aperiodic cases, respectively. In addition, Wang et al. [19] proposed an asymptotically optimal LAZ sequence set based on a cubic function. Consequently, this drives us to develop sequences with zero or low AF sidelobe characteristics in desired regions. The specific parameters of the above mentioned sequences are listed in Table I. In column 6 of Table I, P.O, P.AO, and AP.AO. mean optimal with periodic, asymptotically optimal with periodic, and asymptotically optimal with aperiodic, respectively.
@@
  Ref. Length Set size $Z_{x}$ $Z_{y}$ ${\theta}_{\text{max}}(\hat{\theta}_{\text{max}})$ Optimality   Constraints   [3] q1q-1 11 q1q-1 q1q-1 $\sqrt{q}$ P.O. $q=p^{l}$, $p$ is a prime. [13] pp 11 $p$ $p$ $\sqrt{p}$ P.O. $p$ is an odd prime. [13] $L$ 11 $\frac{L}{r}$ $r$ 0 P.O. $\gcd(a,L)=1\,\textrm{if $L$ is odd },$ $r=\gcd(2a,L)>1$. [13] $L$ NN $\frac{\lfloor L/N\rfloor}{r}$ $r$ 0 P.O, if $N|L$. $\gcd(a,L)=1\,\textrm{if L is odd },$ $r=\gcd(2a,L)>1$. [17] $^{2} MNMN$ $\lfloor\frac{N}{K}\rfloor$ $K$ 0 P.AO. $K<N$, $\gcd(K,N)=1$. [17] $N(KN+P)$ NN NN $K$ 0 P.AO. $\gcd(P,NK)=1$. [17] p(p1)p(p-1) $p$ p1p-1 $p$ $\sqrt{p}$ P.AO. $p$ is an odd prime. [19] NN $M$ $p$ $\lfloor\frac{N}{M}\rfloor$ $\sqrt{N}$ P.AO NN  is an odd number, $p$ is the smallest prime of $N$, $M\leq N$. [18] $^{2} NN$ $p$ NN NN, N+p1N+p-1 P.AO, AP.AO. NN  is an odd number, $p|N(\textrm{the smallest prime})$. [18] NKNK NN $p$ KN+1K-N+1 $K$, K+p1K+p-1 P.AO, AP.AO. NN  is an odd number, N<K<2N1N<K<2N-1, $p|N(\textrm{the smallest prime})$. Theorem4 (i), Theorem6 (i) NKNK NN p1p-1 $K$ $K$, K+p1K+p-1 P.AO, AP.AO. $K=p^{e},\textrm{$p$\,is an odd prime},$ $N=\varphi(p^{e})$. Theorem4 (ii) NKNK NN NN $p$ K+N1K+N-1 P.AO $K=p^{e},\textrm{$p$\,is an odd prime},$ $N=\varphi(p^{e})$. Theorem8 $K_{1} NN$ p1p-1 $K_{1}-K+2$ $K_{1}$, $K_{1}+p-2$ P.AO, AP.AO. $p^{e},\textrm{$p$\,is an odd prime}, KK$$\leq K_{1}<2K$, $N=\varphi(p^{e})$.
In this paper, we show some generic new constructions on LPNFs in Section 3, and by using these new LPNFs, we present a series of LAZ sequence sets with new parameters. Some sequence sets of the proposed are asymptotically optimal with respect to YZFLLT bounds both periodic and aperiodic cases, respectively. Before doing our main results, in Section 2, we introduce some preliminaries on LAZ sequence sets and LPNFs, including some related results we need in this paper.
@@ II LAZ Sequence Sets and Locally Perfect Nonlinear Functions
In this section, we review some fundamental concepts related to LAZ sequence sets and LPNFs, along with several results that will be used in the subsequent analysis.
@@ II-A LAZ Sequence Sets
Throughout the rest of the paper, let $\mathbb{C}$ denote the set of all complex numbers and $\mathcal{S}$ denote the set of sequences of length $D$ which has a size of $M$, all sequences $\mathbf{a}=(a(0),a(1),a(2),\cdots,a(D-1))$ are regarded as unimodular sequences with length $D$. Specifically, ${a}(k)\in\mathbb{C}$ and |a(k)|=1|{a}(k)|=1 for all $0\leq k\leq{D-1}$. In particular, $D$ is the period of $\mathbf{a}$ if a(k+D)=a(k){a}(k+D)={a}(k) for all $k\geq 0$. $\omega_{N}=e^{\frac{2\pi\sqrt{-1}}{N}}=e^{\frac{2\pi i}{N}}$ is a $N$-th primitive complex root of unity.
Let $\mathbf{a}$ and $\mathbf{b}$ be unimodular sequences with length $D$. Then the aperiodic cross-ambiguity function (cross-AF) of $\mathbf{a}$ and $\mathbf{b}$ at time shift $\tau$ and Doppler shift $\nu$ is defined by:
@@
where for $\in\mathbb{C}, c¯\overline{c}$ represents complex conjugate of $c$.
@@
where $t+\tau$ to be taken modulo $D$. We denote $\widehat{AF}_{\mathbf{a}}(\tau,\nu)=\widehat{AF}_{\mathbf{a},\mathbf{a}}(\tau,\nu)$ and ${AF}_{\mathbf{a}}(\tau,\nu)={AF}_{\mathbf{a},\mathbf{a}}(\tau,\nu)$, called the auto-AF of $\mathbf{a}$, aperiodic and periodic, respectively.
@@
Lemma 1 (Ye–Zhou–Fan–Liu–Lei–Tang bounds [13]).
For any unimodular periodic or aperiodic sequence set: $\mathcal{S}=(M,D,\Pi,\theta_{\text{max}})$ or $\mathcal{S}=(M,D,\Pi,\hat{\theta}_{\text{max}})$, where $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$. We have
@@
Further assume that $\rho_{\text{LAZ}}=\frac{\theta_{\text{max}}}{\Delta}(\geq 1)$ and $\hat{\rho}_{\text{LAZ}}=\frac{\hat{\theta}_{\text{max}}}{\hat{\Delta}}(\geq 1)$, $\Delta$ and $\hat{\Delta}$ are YZFLLT bounds, $\rho_{\text{LAZ}}$ and $\hat{\rho}_{\text{LAZ}}$ are conventionally referred to as optimality factors. The optimality factors $\rho_{\text{LAZ}}$ and $\hat{\rho}_{\text{LAZ}}$ serve as metrics to evaluate how closely the constructed sequences approach the theoretical lower bounds in both periodic and aperiodic cases, respectively. i.e.
@@
Remark 1.
When the parameters $M$, $Z_{x}$, and $Z_{y}$ satisfy either $Z_{x}>\sqrt{\frac{3N^{2}}{MZ_{y}}}$ with $MZ_{y}\geq 3$, or $Z_{x}>\frac{\pi}{\gamma}$ with $5\leq MZ_{y}\leq N^{2}$, where $\gamma=\arccos\left(1-\frac{MZ_{y}}{N^{2}}\right)$, a tighter aperiodic AF lower bound was derived by Meng et al. in [14, Corollaries 2 and 3], improving upon the result given in [13, Theorem 4]. Nevertheless, since our constructed LAZ sequence set does not fulfill either of these conditions, we adopt the lower bound for aperiodic AF provided by Ye et al. in [13, Theorem 4].
@@ II-B Locally Perfect Nonlinear Functions
Let $K\geq N$ be two positive integers, and $f:$ $\mathbb{Z}_{N}\rightarrow\mathbb{Z}_{K}$ be a function, where $\mathbb{Z}_{N}=\mathbb{Z}/N\mathbb{Z}$. For $0<Z_{x}\leq N$ and $0<Z_{y}\leq K$, The function $f$ is called a locally perfect nonlinear function and denoted by $\langle N,K,Z_{x},Z_{y}\rangle$-LPNF if for any $Z_{y}<b<Z_{y}$ and $Z_{x}<a<Z_{x}$, $a\neq 0$, the equation f(x+a)f(x)=bf(x+a)-f(x)=b has at most one solution $x\in\mathbb{Z}_{N}$.
@@
If $Z_{x}\geq\left\lceil\frac{N+1}{2}\right\rceil$, then the interval $Z_{x}<a<Z_{x}$ covers all elements $a\in\mathbb{Z}_{N}$. Similarly, if $Z_{y}\geq\left\lceil\frac{K+1}{2}\right\rceil$, then $Z_{y}<b<Z_{y}$ covers all elements $b\in\mathbb{Z}_{K}$.
The notion of LPNF was introduced in [18] and successfully used to construct LAZ sequence sets by the interleaving technique. The result is as follows.
Lemma 2 ([18] Theorems 1 and 2).
From an $\langle N,K,Z_{x},Z_{y}\rangle$-LPNF, a LAZ sequence set can be constructed with parameters $(M,D,\Pi,\theta_{\max})$ or $(M,D,\Pi,\hat{\theta}_{\max})$, where $M=N$, $D=NK$, $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$, $\theta_{\max}=K$, and $\hat{\theta}_{\max}=K+Z_{x}-1$.
Wang et al.[18] gave a series of LPNFs and constructed a series of LAZ sequence sets, which are presented in the above theorem. Some of them are asymptotically optimal in both periodic and aperiodic cases. In the next section, we will reveal some new constructions of LPNFs, which provides significant improvements over existing frameworks [18]. And then produce a series of LAZ sequence sets with new parameters that are not covered by the parameters of [18]. Furthermore, some classes of the proposed sequence sets demonstrates an asymptotically optimal property in both periodic and aperiodic cases.
III Proposed Constructions of LPNFs
For $K\geq 3$, let $\mathbb{Z}_{K}^{\ast}$ denote the multiplicative group of all invertible elements of the ring $\mathbb{Z}_{K}$. Then $\mathbb{Z}_{K}^{\ast}$ contains $\varphi(K)$ elements, where $\varphi(\cdot)$ is the Euler’s total function. In this section, our primary objective is to construct three new classes of LPNFs. The following lemma provides a necessary foundation for the subsequent constructions of LPNFs.